OpenStudy (anonymous):

When evaluating the following expression, which should you do first? 17- 15 + 5

OpenStudy (anonymous):

-15+5

OpenStudy (anonymous):

addition and subtraction are commutative, the order does not matter, In any case you could do it in the order it is written

OpenStudy (anonymous):

proceed from left to right 17 - 5 = 2 2 + 5 = 7

OpenStudy (anonymous):

addition is commutative but not subtraction

OpenStudy (anonymous):

+a-b=-b+a ur right being strict I am adding